OpenStudy (anonymous):

simplify (1/x+h)-(1/h)

OpenStudy (anonymous):

\[ \frac 1{x+h} - \frac 1 h=\frac {h-(x+h)}{h(x+h)} \] Can you finish it?

OpenStudy (anonymous):

\[ \frac 1{x+h} - \frac 1 h=\frac {h-(x+h)}{h(x+h)}=\frac{-x}{h(x+h)} \]

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anonymous):

yw